Each blade of grass is busy undergoing photosynthesis in the following UNBALANCED equation: CO2 + H2O + light --> C6H12O6 + O2. How many liters of oxygen gas could a plant produce from 125.67 g of water?
1
Expert's answer
2020-10-16T08:24:19-0400

At first, the equation should be balanced:

6CO2+6H2O+light=C6H12O6+6O2.

Volume of oxygen may be found from a proportion:

V(O2)=m(H2O)×6×Vm/6×M(H2O)=m(H2O)×Vm/M(H2O),

where Vm is a molar volume and M(H2O) is a molar mass of water.

V(O2)=125.67×22.4/18=156.39(L)